# SegSTRONG-C: Segmenting Surgical Tools Robustly On Non-adversarial Generated Corruptions – An EndoVis’24 Challenge

### v1

@@ Abstract
The dataset consists of mock endoscopic video sequences with corresponding binary segmentation masks for the robot tool segmentation task. We mock the endoscopic scene with two patient-side manipulators (PSMs) and animal tissue backgrounds to ensure a photo-realistic appearance. We manually teleoperate the robot to generate the trajectory in free space. The binary segmentation masks serve as ground truth annotations for surgical tools. The dataset consists of $17$ sequences, collected from different robot and camera configurations and different robot trajectories. Each sequence is collected at $10$ frames per second and consists of $300$ uncorrupted frames for the left camera and $300$ uncorrupted frames for the right camera. We collect all corrupted versions (background, smoke, bleeding, and low brightness) for each sequence under the same configuration and trajectory. We provide $11$ sequences of the dataset with only uncorrupted frames as the train set. We provide $3$ video sequences with uncorrupted frames and corresponding frames with background corruption as the validation set. We use the final 3 sequences with certain corruptions(smoke, bleeding, and low brightness) as the test set. The models submitted by the challenge participants will be tested on this test set consisting of sequences with photo-realistic non-adversarial corruptions. The models will be tested on each corruption separately. As the challenge is currently active, we withhold the release of the test set. Table 1 shows the summary of the dataset. All corrupted versions for the training and validation will be released after the challenge submission. The example images are shown in Figure 1.
@@
Test set 9 15 - - 300 + 300 300 + 300 300 + 300
9 16 - - 300 + 300 300 + 300 300 + 300
9 17 - - 300 + 300 300 + 300 300 + 300
Figure 1: Example images for regular image and non-adversarial corruptions(background, smoke, bleeding, and low brightness.).
@@
The data was acquired in the Robotorium of the Laboratory for Computational Sensing and Robotics (LCSR), Johns Hopkins University by surgical robotics experts familiar with operating the da Vinci robot via dVRK [18]. We use dVRK [18] as our robot platform with the endoscopic camera manufactured by SCHÖLLY and the image process unit manufactured by Ikegami as the perception part. We first adjust the camera/robot configuration and perform calibrations. For each configuration, we collect different trajectories by human operation. For each trajectory, we replay the recorded kinematics to reproduce the trajectory and record video sequences under different scenarios including non-corrupted and corrupted versions with background, smoke, bleeding, and low brightness. To generate non-adversarial corruptions for the same testing samples, we record the robot trajectory, replay the kinematics via dVRK, and manually add corruption for each replay. The background corruption is generated by changing the type of background tissue. The smoke corruption is mimicked by adding artificial fog via a fog machine. The bleeding corruption is mimicked by fake blood. The low brightness corruption is generated by turning down the camera light. The data collection pipeline is shown in the following steps:
@@
The annotations are generated via a semi-automatic pipeline based on exclusively collected sequences with purely black backgrounds where the only salient areas in the images are the PSMs. We first automatically generate a segmentation mask for the PSMs via the segment anything model (SAM) [36] where the prompts are generated via traditional background extraction algorithm. We select experts who are familiar with the robot and segmentation task to be the annotators to verify and correct the automatically generated annotation. The annotation pipeline can be expressed in the following steps:
  • Step 1: Prompt generation - We use a traditional background extraction algorithm to generate rough masks for the foreground PSMs. The rough masks are converted to the prompt points and bounding boxes for SAM.
@@
  • Step 3: Failure case selection - Human annotators are involved in examining the generated mask from SAM and selecting the failure cases. We had two annotators examining the same sequences and using the union of their selections as the set of failure cases.
@@
Metrics:
We use dice similarity coefficient (DSC) and normalized surface distance (NSD) averaged from different tolerances for the robot tool for multiple corruptions - low brightness, smoke, and blood. The DSC is a widely used metric in the field of medical image analysis. For a 2D binary class image segmentation task, DSC is defined as
@@
, where $TP$ is the number of true positive pixels for the class, $FP$ is the number of false positives and $FN$ is the number of false negatives. NSD measures the overlap of two boundaries between the predicted and ground truth segmentation masks. A boundary pixel is counted as overlapping when the closest distance to the other boundary is less or equal to the specified tolerance. NSD for a 2D binary class image segmentation task is defined as:
@@
, where $A$ and $B$ are the masks, $S_{A}$ and $S_{B}$ are boundaries, and $\mathcal{B}_{A}^{(\mathrm{\tau})}$ and $\mathcal{B}_{B}^{(\mathrm{\tau})}$ are the boundary of $A$ and $B$ with an extended border of width tolerance $\tau$. We calculate the average DSC and NSD over images. DSC is the standard metric for segmentation while NSD is complementary to DSC. DSC reveals performance more on the chunk area while NDS focuses on the boundary. We evaluate the performance only on the unseen domain to encourage participants to focus on the algorithm’s robustness.
Ranking:
@@
We rank algorithms with total scores summed over 3 test domains - bleeding, smoke, and low brightness. Each Domain will have two ranks based on the DSC and NSD metrics. For ranking each domain, we will first sort the method based on the metrics over images. Then, we do a pairwise significance test for adjacent methods to test whether they have statistically significant differences in performance. If yes, we rank them according to the order, otherwise they have the same rank. The mean value indicates the overall performance, and the pairwise significance test can justify whether one method has a significant advantage over the other one. We give points in a decreasing arithmetic sequence according to the rank for each test domain. For example, Rank from 1 to 5 and get points from 5 to 1 if there are 5 participants in total. We only test on the unseen corrupted domains as we care about the overall performance of the model on the corrupted domain. Since we have more than one test domain, the summation of points given by rank is a reasonable way to have fair importance for each domain. Using rank and ignoring absolute performance difference is to avoid the result from one specific domain and metric dominating others. Since the goal for the participating team is to develop methods that are robust against all corruptions instead of algorithms achieving dominating performance on one domain but failing to generalize to others, The proposed overall ranking method better suits the goal of the challenge.
Baseline Analysis
@@
In this section, we first briefly introduce the selected segmentation networks and data augmentation methods. Then, we present the testing results of these methods and perform a brief analysis of the baseline results.
Networks
U-Net
: U-Net[2] is originally developed for biomedical image segmentation. The architecture of U-Net consists of two paths: a contracting path and an expansive path. The contracting path is characterized by repeated convolutions and resembles a conventional convolutional network, followed by a rectified linear unit(RELU) and a max pooling operation. On the other hand, the expansive path enhances feature and spatial information through a series of up-convolutions and concatenation of features from the contracting path. Our implementation uses a 4-level vanilla U-Net as the baseline model.
@@
: The most recent version of the DeepLab family is DeepLabV3+ [3]. The architecture is changed to an encoder-decoder structure. The encoder part makes use of atrous separable convolution which separates the convolution process into a Depthwise Convolution and a Pointwise Convolution. Depthwise Convolution applies different kernels to each channel, and Pointwise Convolution combines the outputs across channels. The decoder part combines the low-level features and the upsampled output from the encoder to recover the spatial information by upscaling.
@@
: SegFormer [7] is a semantic segmentation framework that combines a hierarchical transformer encoder with a multi-layer perceptron (MLP) decoder. The hierarchically structured transformer encoder avoids the interpolation of positional codes and outputs multi-scaled features. The lightweight MLP decoder combines both local and global attention to aggregate the multi-layer information.
@@
: SETR [6] provides an alternative perspective to the encoder-decoder-based FCN architecture. It treats semantic segmentation as a sequence-to-sequence prediction task in which the author splits the image into patches, linearly projects each patch, adds positional embeddings, and feeds the sequence into the transformer encoder. Three different decoders are designed to perform pixel-level segmentation, and we implement all three: (1) Naive upsampling which projects the transformer feature to the dimension of class number and directly upsample to the desired image resolution, (2) Progressive upsampling which alternates convolution and upsampling operations instead of one-step upscaling, and (3) Multi-Level feature aggregation which takes feature representations from $M$ transformer layers uniformly, reshapes and aggregates the feature map top-down, fuses features from all channels and upsamples to the full resolution.
Augmentation
Figure 2: Example image augmentations
AutoAugment
@@
: AutoAugment [30] searches for the best sequence of transformations that improves the performance of a model on a given dataset. We modify the Pytorch implementation of the AutoAugmentPolicy class and choose the policy learned on ImageNet.
Elastic
: Elastic transformation generates displacement vectors for each image pixel based on random offsets. It will transform the image given $\alpha$ and $\beta$, where $\alpha$ controls the strength and $\beta$ controls the smoothness of displacements. We modify the Pytorch implementation of the ElasticTransform class. Let $\alpha_{x}$ and $\alpha_{y}$ be the magnitude of displacements, $\sigma_{x}$ and $\sigma_{y}$ be the smoothness of displacements along $x$ and $y$ axis, and $h$ and $w$ be the height and width of the image. Kernel size along $x$ axis is: $k_{x}=\text{int}(8\sigma_{x}+1)$ thus half of the kernel size is $k_{x\text{\_half}}=\frac{1}{2}(k_{x}-1)$, and the displacement vector $d_{x}$ is first initialized as an array of random numbers between $-1$ and $1$. Each pixel is defined as the Gaussian blur of $d_{x}$ times $\alpha_{x}$ and divided by image height $h$. The 2D Gaussian kernel $h_{x}$ is defined as $h_{x}[m,n]=h_{1}[m]h_{1}[n]^{T}$ where $h_{1}$ is the normalized discrete 1D Gaussian kernel of length $k_{x}$. $d_{x}$, $d_{y}$ is defined as $d_{x}[m,n]=\frac{\alpha_{x}}{h}d_{x}[m,n]*h_{x}[m,n]$, $d_{y}[m,n]=\frac{\alpha_{y}}{w}d_{y}[m,n]*h_{y}[m,n]$. The complete displacement vector field is obtained by concatenating $d_{x}$ and $d_{y}$. The detailed derivations of $h_{1}$, $d_{x}$, and $d_{y}$ are in the equation:
$h_{1}[m]=\frac{\exp[(-\frac{1}{2}(\frac{-k_{x\text{\_half}}+m}{\sigma_{x}})^{2})]}{\sum_{n=-k_{x\text{\_half}}}^{k_{x\text{\_half}}}\exp[(-\frac{1}{2}(\frac{-k_{x\text{\_half}}+n}{\sigma_{x}})^{2})]}$
$d_{x}[m,n]=\frac{\alpha_{x}}{h}\sum_{j=-\infty}^{\infty}\sum_{i=-\infty}^{\infty}d_{x}[i,j]\cdot h_{x}[m-i,n-j]$
$d_{y}[m,n]=\frac{\alpha_{y}}{w}\sum_{j=-\infty}^{\infty}\sum_{i=-\infty}^{\infty}d_{y}[i,j]\cdot h_{y}[m-i,n-j]$
Projective
: Projective transformation transforms the image such that it is viewed from a different viewing perspective with a given probability. We modify the Pytorch implementation of the RandomPerspective class. The four new coordinates of the corners of the image are first randomly generated according to the input distortion scale. In perspective transform, each pixel $(x,y)$ in the original image is mapped to its new coordinate by the formula: $(u_{i},v_{i})=(\frac{ax_{i}+by_{i}+c}{gx_{i}+hy_{i}+1},\frac{dx_{i}+ey_{i}+f}{gx_{i}+hy_{i}+1})$ for $i=0,1,2,3$. In matrix form, it can be written as:
$\begin{bmatrix}x_{0}&y_{0}&1&0&0&0&-x_{0}u_{0}-y_{0}u_{0}\\ x_{1}&y_{1}&1&0&0&0&-x_{1}u_{1}-y_{1}u_{1}\\ x_{2}&y_{2}&1&0&0&0&-x_{2}u_{2}-y_{2}u_{2}\\ x_{3}&y_{3}&1&0&0&0&-x_{3}u_{3}-y_{3}u_{3}\\ 0&0&0&x_{0}&y_{0}&1&-x_{0}v_{0}-y_{0}v_{0}\\ 0&0&0&x_{1}&y_{1}&1&-x_{1}v_{1}-y_{1}v_{1}\\ 0&0&0&x_{2}&y_{2}&1&-x_{2}v_{2}-y_{2}v_{2}\\ 0&0&0&x_{3}&y_{3}&1&-x_{3}v_{3}-y_{3}v_{3}\end{bmatrix}\cdot\begin{bmatrix}a\\ b\\ c\\ d\\ e\\ f\\ g\\ h\end{bmatrix}=\begin{bmatrix}u_{0}\\ u_{1}\\ u_{2}\\ u_{3}\\ v_{0}\\ v_{1}\\ v_{2}\\ v_{3}\end{bmatrix}$
Solve this system to obtain the least-square solution for the coefficients, and map each pixel to its new position by the perspective transformation formula.
Results
We evaluate the performance of baseline models with and without data augmentation across different domains. While all models show reasonable accuracy under regular conditions, their effectiveness substantially deteriorates when tested in shifted domains such as in the presence of bleeding, low brightness, and smoke. The data augmentations applied, including AutoAugment, elastic, and projective transformations, generally failed to improve model performance in these domains. Notably, UNet combined with AutoAugment showed a marginal improvement under smoke conditions but this improvement is inconsistent across all domains. As shown in Figure 3, under regular conditions, the models adequately segmented tools from the background, but struggled in areas with appearance change and subtle contrast changes. Failures are predominantly under bleeding conditions due to the appearance (color) change of the shaft, leading to incorrect segmentation of the shaft. Under low brightness conditions, the darker the region is, the more likely a failure in segmentation will occur. The pervasive false positives under smoke conditions indicate that the models interpret the bright regions (smoke) as the foreground. These results suggest that all baseline models are highly dependent on adequate targets’ saliency to successfully distinguish between the tools and the background. They are also sensitive to any appearance change for the tool (bleeding) and background (smoke).
Models NSD DSC
Bleeding Smoke Low Brightness Bleeding Smoke Low Brightness
DeepLabv3 [3] 0.5629 0.4637 0.4000 0.6896 0.6538 0.5352
Segformer [7] 0.5133 0.5266 0.4194 0.6802 0.6906 0.6145
SETR_PUP [6] 0.2531 0.3354 0.2599 0.4717 0.5848 0.4053
SETR_MLA [6] 0.2798 0.3374 0.1571 0.5358 0.6206 0.3385
SETR_Naive [6] 0.3312 0.4409 0.0092 0.5064 0.6657 0.0275
UNet [2] 0.5677 0.5084 0.4390 0.7052 0.6603 0.5750
UNet [2] + AutoAugment [30] 0.6654 0.8152 0.5344 0.7910 0.8895 0.6965
UNet [2] + Elastic 0.5622 0.5207 0.3931 0.6910 0.6583 0.5190
UNet [2] + Projective 0.5661 0.5702 0.4265 0.6903 0.6978 0.5792
SETR [6] + AutoAugment [30] 0.1172 0.0910 0.0773 0.1918 0.2157 0.0934
SETR [6] + Elastic 0.2836 0.3363 0.1654 0.5350 0.6192 0.3472
SETR [6]+ Projective 0.2240 0.3055 0.0552 0.4168 0.5766 0.1447
Conclusion
We present the SegSTRONG-C challenge, aiming to assess algorithms’ robustness to unforeseen yet plausible complications in surgical scenarios. To generate non-adversarial corruptions we use and refine the data collection and annotation pipeline from CaRTS [16] and collect 17 sequences for training and evaluation. We use DSC and NSD as quantitative metrics to evaluate algorithms’ performance on the testing set and propose a balanced ranking method to alleviate the influence of extreme performance on a single domain while emphasizing the statistical significance during ranking. The baseline analysis provides baseline performance for participants to refer to while providing a basic insight that the models’ success relies on adequate saliency and that all models are sensitive to appearance change.
Acknowledgement:

### v2

@@ 5.1 Dataset
The dataset consists of mock endoscopic video sequences with corresponding binary segmentation masks for the robot tool segmentation task. We mock the endoscopic scene with two patient-side manipulators (PSMs) from the da Vinci surgical robot and animal tissue backgrounds to ensure a photo-realistic appearance. We manually teleoperate the robot to generate the trajectory in free space. The binary segmentation masks serve as ground truth annotations for surgical tools. The dataset consists of $17$ sequences, collected from different robot and camera configurations and different robot trajectories. Each sequence is collected at $10$ frames per second and consists of $300$ uncorrupted frames for the left camera and $300$ uncorrupted frames for the right camera. We collect all corrupted versions (background, smoke, bleeding, and low brightness) for each sequence under the same configuration and trajectory. We provide $11$ sequences of the dataset with only uncorrupted frames as the train set. We provide $3$ video sequences with uncorrupted frames and corresponding frames with background corruption as the validation set. We use the final 3 sequences with certain corruptions(smoke, bleeding, and low brightness) as the test set. The models submitted by the challenge participants are tested on this test set consisting of sequences with photo-realistic non-adversarial corruptions. The models are tested on each corruption separately. Table 5 shows the summary of the dataset. All corrupted versions for the training and validation are released and can be found on the challenge website segstrongc.cs.jhu.edu. The example images are shown in Figure 4.
@@
Test set 9 15 300 + 300 300 + 300 300 + 300 300 + 300 300 + 300
9 16 300 + 300 300 + 300 300 + 300 300 + 300 300 + 300
9 17 300 + 300 300 + 300 300 + 300 300 + 300 300 + 300
Figure 4: Example images for uncorrupted image and non-adversarial corruptions(background, smoke, bleeding, and low brightness.).
@@
The data was acquired in the Robotorium of the Laboratory for Computational Sensing and Robotics (LCSR), Johns Hopkins University, by surgical robotics experts familiar with operating the da Vinci robot via dVRK [43]. We use dVRK [43] as our robot platform with the endoscopic camera manufactured by SCHÖLLY and the image process unit manufactured by Ikegami as the perception part. We refer to the data collection pipeline in CaRTS [40, 41] to generate paired corrupted video sequences. We first adjust the camera/robot configuration and perform calibrations. For each configuration, we collect different trajectories by human operation. For each trajectory, we replay the recorded kinematics to reproduce the trajectory and record video sequences under different scenarios, including non-corrupted and corrupted versions with background, smoke, bleeding, and low brightness. To generate non-adversarial corruptions for the same testing samples, we record the robot trajectory, replay the kinematics via dVRK, and manually add corruption for each replay. The background corruption is generated by changing the type of background tissue. The smoke corruption is mimicked by adding artificial fog via a fog machine. The bleeding corruption is mimicked by fake blood. The low brightness corruption is generated by turning down the camera light. The data collection pipeline is shown in the following steps:
@@
The annotations are generated via a semi-automatic pipeline based on exclusively collected sequences with purely black backgrounds, where the only salient areas in the images are the PSMs. We first automatically generate a segmentation mask for the PSMs via the segment anything model (SAM) [44], where the prompts are generated via a traditional background extraction algorithm [45]. We select experts who are familiar with the da Vinci robot and segmentation task to be the annotators to verify and correct the automatically generated annotation. The annotation pipeline can be expressed in the following steps:
  • Step 1: Prompt generation - We use a traditional background extraction algorithm to generate coarse masks for the foreground PSMs. The coarse masks are converted to the prompt points and bounding boxes for SAM.
@@
  • Step 3: Failure case selection - Human annotators are involved in examining the mask generated from SAM and selecting the failure cases. We had two annotators examining the same sequences and using the union of their selections as the set of failure cases.
@@ 5.2 Evaluation
Metrics
We use the DICE similarity coefficient (DSC) and normalized surface distance (NSD) averaged from different tolerances for the robot tool for multiple corruptions - low brightness, smoke, and blood. The DSC is a widely used metric in the field of medical image analysis. For a 2D binary class image segmentation task, DSC is defined as
@@
, where $TP$ is the number of true positive pixels for the class, $FP$ is the number of false positives, and $FN$ is the number of false negatives. NSD measures the overlap of two boundaries between the predicted and ground truth segmentation masks. A boundary pixel is counted as overlapping when the closest distance to the other boundary is less than or equal to the specified tolerance. NSD for a 2D binary class image segmentation task is defined as:
@@
, where $A$ and $B$ are the masks, $S_{A}$ and $S_{B}$ are boundaries, and $\mathcal{B}_{A}^{(\mathrm{\tau})}$ and $\mathcal{B}_{B}^{(\mathrm{\tau})}$ are the boundary of $A$ and $B$ with an extended border of width tolerance $\tau$. We calculate the average DSC and NSD over images. DSC is the standard metric for segmentation, while NSD is complementary to DSC. DSC reveals performance more on the chunk area while NDS focuses on the boundary. We evaluate the performance only on the unseen domain to encourage participants to focus on the algorithm’s robustness.
Ranking
@@
We rank algorithms with total scores summed over 3 test domains - bleeding, smoke, and low brightness. Each domain will have two ranks based on the DSC and NSD metrics. For ranking each domain, we will first sort the method based on the metrics over images. Then, we do a pairwise significance test for adjacent methods to test whether they have statistically significant differences in performance. If yes, we rank them according to the order, otherwise, they have the same rank. The mean value indicates the overall performance, and the pairwise significance test can justify whether one method has a significant advantage over the other one. We give points in a decreasing arithmetic sequence according to the rank for each test domain. For example, Rank from 1 to 5 and get points from 5 to 1 if there are 5 participants in total. We only test on the unseen corrupted domains as we care about the overall performance of the model on the corrupted domain. Since we have more than one test domain, the summation of points given by rank is a reasonable way to have fair importance for each domain. Using rank and ignoring absolute performance difference is to avoid the result from one specific domain and metric dominating others. Since the goal for the participating team is to develop methods that are robust against all corruptions instead of algorithms achieving dominating performance on one domain but failing to generalize to others, the proposed overall ranking method better suits the goal of the challenge.
5.3 Challenge Logistics
SegSTRONG-C challenge launches as a one-time sub-challenge under the EndoVis2024 challenge [^1] https://opencas.dkfz.de/endovis/challenges/2024/. The challenge launches on May 1st, 2024, with all training and validation datasets released. All submissions were collected before September 6th, 2024, and evaluated using docker containers. We announced one winner and two runner-ups on MICCAI2024 (October 10th, 2024). The final results were released on October 12th, 2024, and test data were released on October 18th, 2024.
Registration and Submissions
In total, we received 40 registrations from all over the world. All related countries are listed below: Canada (12.5%), USA (7.5%), Republic of Korea (10%), China (32.5%), Switzerland(2.5%), Japan(2.5%), Portugal(2.5%), India(10%), Germany(7.5%), UK(2.5%), Iran(2.5%), Denmark(2.5%), Singapore(2.5%), Morocco(2.5%). Within all registrations, we received 10 valid submissions. They are from Canada (20%), USA (20%), Republic of Korea (20%), China (10%), Switzerland(10%), Japan(10%), Portugal(10%).
5.4 Baseline Models
@@
In this section, we introduce the selected segmentation networks and data augmentation methods for baseline analysis.
@@
The most recent version of the DeepLab family is DeepLabV3+ [17]. The architecture is changed to an encoder-decoder structure. The encoder part makes use of atrous separable convolution which separates the convolution process into a Depthwise Convolution and a Pointwise Convolution. Depthwise Convolution applies different kernels to each channel, and Pointwise Convolution combines the outputs across channels. The decoder part combines the low-level features and the upsampled output from the encoder to recover the spatial information by upscaling.
@@
SegFormer [21] is a semantic segmentation framework that combines a hierarchical transformer encoder with a multi-layer perceptron (MLP) decoder. The hierarchically structured transformer encoder avoids the interpolation of positional codes and outputs multi-scaled features. The lightweight MLP decoder combines both local and global attention to aggregate the multi-layer information.
@@
SETR [20] provides an alternative perspective to the encoder-decoder-based FCN architecture. It treats semantic segmentation as a sequence-to-sequence prediction task in which the author splits the image into patches, linearly projects each patch, adds positional embeddings, and feeds the sequence into the transformer encoder. Three different decoders are designed to perform pixel-level segmentation, and we implement all three: (1) Naive upsampling which projects the transformer feature to the dimension of class number and directly upsample to the desired image resolution, (2) Progressive upsampling which alternates convolution and upsampling operations instead of one-step upscaling, and (3) Multi-Level feature aggregation which takes feature representations from $M$ transformer layers uniformly, reshapes and aggregates the feature map top-down, fuses features from all channels and upsamples to the full resolution. We apply naive upsampling in our baseline analysis.
UNet
UNet[16] was originally developed for biomedical image segmentation. The architecture of UNet consists of two paths: a contracting path and an expansive path. The contracting path is characterized by repeated convolutions and resembles a conventional CNN, followed by a rectified linear unit(RELU) and a max pooling operation. On the other hand, the expansive path enhances feature and spatial information through a series of up-convolutions and concatenation of features from the contracting path. Our implementation uses a 4-level vanilla UNet as the baseline model.
UNet + AutoAugment
@@
AutoAugment [34] searches for the best sequence of transformations that improves the performance of a model on a given dataset. We modify the Pytorch implementation of the AutoAugmentPolicy class and choose the policy learned on ImageNet. We apply AutoAugment during the training of UNet for this baseline.
SAM2
SAM2 [35] extends the Segment Anything Model (SAM) [44] for video segmentation by incorporating temporal information. It retains SAM’s core components—image encoder, prompt encoder, and mask decoder—while introducing a memory encoder and memory bank to capture frame-to-frame consistency. The memory mechanism allows SAM2 to leverage past predictions, enhancing temporal coherence in segmentation. We adopt SAM2 as a baseline for its strong performance in point prompt-based video segmentation and its ability to model temporal context effectively. We use the ground truth mask to sample 5 positive points and 5 negative points for the first frame of each sequence.
5.5 Challenge Submissions
Figure 5: Overview of our proposed corruption simulation + segmentation framework. (Team VSI)
VSI (Yicheng Leng, Seok Bong Yoo, Eung-Joo Lee):
Team VSI’s framework is illustrated in Figure 5. They use PSPNet50  [46] for the surgical tool segmentation. To improve the model’s robustness against different types of corruption, they apply image augmentation techniques from ImageNet-C [31] to simulate the smoke and low-brightness cases. They ignore the augmentation of the bleeding case as they think the color and shape of red artifacts added onto the image are difficult to determine, and the bleeding area brings minimal pixel-wise impact on both the surgical tool and tissue background.
UBCRCL (Zijian Wu, Yuxin Chen, Septimiu E. Salcudean):
RGB-D semantic segmentation methods have been demonstrated to improve the segmentation performance by fusing depth information [47]. Team UBCRCL tries to explore the performance of an RGB-D method on surgical datasets, especially when confronting corrupted images. They adopt a state-of-the-art method, DFormer [48], a pretraining framework to encode RGB-D representations. They apply an off-the-shelf DepthAnything [49] model to generate pseudo depth for each frame, as the depths from stereo-matching [50] are not satisfying without rectified images. To enhance the model’s robustness, they perform the 3D Common Corruption (3DCC) [51] for data augmentation. Specifically, they tune the following augmentation to different degrees during the training: reducing the image brightness, the color and density of fog, and random occlusions.
SAM_KHU(Samra Irshad, Shadi Albarqouni, Seong Tae Kim):
Figure 6: (a). VM-UNet architecture. (b) Details of components in VSS block (Team SAM_KHU)
State space models (SSMs) [52, 53] have recently gained attention for their ability to model long-distance dependencies with linear complexity, improving sequence modeling and computational efficiency, and are now being integrated with architectures like CNNs for tasks such as medical image segmentation [54, 55]. Team SAM_KHU applies Vision Mamba-UNet (VM-UNet), a recently proposed SSM-based model for this challenge. The architecture is shown in Figure 6. The encoder includes VSS blocks from VMamba [56] for feature extraction and patch merging operations for downsampling. The decoder employs VSS blocks and patch-expanding operations to recover the original size of the segmentation output. The skip connection module uses a simple additive operation. To make the model robust to procedural artifacts, they use StarGanv2 [57], an image-to-image translation model that learns the mapping between source and target domains to synthesize these patterns.
Medical_Mechatronics(Yueyi Sun, An Wang, Long Bai, Hongliang Ren):
Figure 7: General workflow of our methodology. (Team Medical_Mechatronics)
For Team Medical_Mechatronics, their primary objective is to introduce comprehensive perturbations onto the pristine images, effectively synthesizing a new, realistic dataset with characteristics mimicking smoke, bleeding, and low-light conditions. By training on this augmented dataset, they aim to develop a more robust model that can generalize well to unseen test data. They have created different sets of datasets for each corruption and trained type-specific models. For smoke, they leverage the widely used fog synthesis approach from the robustness benchmark [31] to generate corruption at different severity levels. For blood, they employ the implementation from the semi-synthetic approach [58], utilizing the BloodDroplet class to introduce blood droplets onto images. For low brightness, they utilize Photoshop’s batch processing capabilities to enhance images and achieve a similar effect. They leverage the great generalization ability of pre-trained visual foundation models in their design. Specifically, they employ Yolo-World [59] to generate bounding box prompts for Segment Anything Model 2 (SAM 2) [35], which will be fine-tuned with their augmented datasets. The overall pipeline of their method is demonstrated in Figure 7.
CCG_DGIST(Ihsan Ullah, Ho-Gun Ha, Attaullah Khan, Hyunki Lee):
Figure 8: The overall architecture of the proposed surgical tool segmentation model. (Team CCG_DGIST)
Team CCG_DGIST utilizes a multi-attention and lightweight model that achieves competitive performance in surgical tool segmentation. The overall framework is shown in Figure 8. This framework incorporates several attention modules to enhance performance: (1) The Dilated Gated Attention Block (DGA), which enables the model to focus more precisely on the target regions by capturing both global and local information. (2) The Inverted External Attention Block (IEA)[60, 61], an efficient external attention mechanism that strengthens the information associations between samples and extracts overarching characteristics from the entire dataset. (3) To capture multistage and multiscale information, they introduce attention bridge modules for both channel and spatial dimensions, termed the Channel Attention Bridge Block (CAB) and Spatial Attention Bridge Block (SAB)[62], which generate their respective attention maps. Additionally, depthwise separable convolutions [63] are employed to reduce the number of parameters and computational complexity, making the model more efficient.
SK(Satoshi Kondo, Satoshi Kasai, Kousuke Hirasawa)
Figure 9: Training de-artifact network and segmentation network. (Team SK)
Team SK applies a two-step regime for the challenge. They use two networks: a de-artifact network and a segmentation network. The left subfigure of Figure 9 shows the de-artifact network. They add one of the three artifacts: brightness, smoke, and blood to the input image for training. The brightness artifacts are realized by reducing the brightness and contrast in a predefined range. The smoke artifacts are realized by alpha-blending an input image and an artificial smoke image. The blood artifacts are realized by adding red-colored ellipses to the input image. The de-artifact network is an encoder-decoder architecture like UNet [16]. The encoder part is replaced by ConvNeXt-v2-base [64]. The loss function is the L1 loss between the input image and the artifact-reduced image. The right subfigure of Figure 9 shows the segmentation network. The artifact addition block has the same functionality as the one when the de-artifact network is trained. For robustness, artifact addition is applied to randomly selected 75 % of the training datasets. The de-artifact network trained in the first step is frozen, and only the segmentation network is trained. The architecture of the segmentation network is the same as the de-artifact network. The loss function is a weighted summation of Dice loss, cross-entropy loss, and Hausdorff distance loss.
Tailor(Sita Tailor, Ricardo Sanchez-Matilla, Imanol Luengo):
Figure 10: Overview of the model and training. (Team Tailor)
Team Tailor applies both surgical-specific and general augmentations to better simulate real surgical conditions. The surgical-specific augmentations included motion blur, low lighting, blood, and smoke, mimicking non-adversarial corruptions typical in operating environments. Additionally, general augmentations like affine transformations, flips, and random resized cropping are applied to introduce further variability. All augmentations were implemented using functions from the Albumentations library [65]. They focus on transformer-based models for robust surgical tool segmentation, selecting the SwinV2 Transformer (SwinV2 small) [66] for its superior performance with segmentation tasks. They modify the model by replacing the classification head with a segmentation head, incorporating a 1x1 convolutional layer followed by upsampling to generate segmentation maps. An overview of the framework is displayed in Figure 10.
FightTumor(Tianhao Fu, Jun Ma, Bo Wang)
Team FightTumor’s approach uses pre-trained UNet [16] with AutoAugment [34]. They customize augmentations in Albumentations [65]. For smoke, the randomized smoke image is merged on top of the original image with an opacity of 1. They also generate a circle flare in which the border is blurred.
ICVS-2AI(Marcos Fernández-Rodríguez, Estevão Lima, João L. Vilaça):
For Team ICVS-2AI, a CycleGAN is trained with default parameters [67] and unpaired images, chosen from a private laparoscopic dataset, with 1496 clear and 220 containing smoke artifacts. To learn the “clear to smoke” style, they use the trained model with the SegSTRONG-C dataset, generate smoke versions of the original images, and use the original ground truth for the new images. Finally, the nnUNet network is trained using standard parameters described in their paper [68], with the increased dataset, using a combined dice and top-k loss.
AIMI(Negin Ghamsarian, Klaus Schoeffmann, Raphael Sznitman)
Team AIMI explore three architectures ReCalNet [69], AdaptNet [70], and DeepPyramid+ [71, 72]. After thorough experiments, they finally apply DeepPyramid+ [71], which features two main modules: Pyramid View Fusion (PVF) and Deformable Pyramid Reception (DPR) as their architecture. The PVF module enhances pixel-level information representation by simulating the human visual system, offering a global-to-local perspective to capture relevant details around each pixel. The DPR module employs dilated deformable convolutions to facilitate shape- and scale-adaptive feature extraction, thereby improving segmentation accuracy for diverse and deformable classes within medical images. They apply common data augmentations like random cropping, color jittering, grayscale transformation, and Gaussian blur during training for better robustness.
Data availability:
The challenge data is released publicly and can be accessed at https://segstrongc.cs.jhu.edu/
Code availability:
The code is made available at https://github.com/hding2455/CaRTS/.
Acknowledgement